OpenStudy (australopithecus):

How do I answer this: A point on a rotating turntable 20.0 cm from the center accelerates from rest to a final speed of 0.700 m/s in 1.75 s. At t = 1.25 s, find the magnitude and direction of (a) the radial acceleration, (b) the tangential acceleration, and (c) the total acceleration of the point.

OpenStudy (anonymous):

angular velocity=0.7 T=1.75 radius=20cm apply formulas for individual questions

OpenStudy (australopithecus):

can you provide me with the formulas?

OpenStudy (anonymous):

\[a= \sqrt{at ^{2}+acp ^{2}}\]

OpenStudy (australopithecus):

is acp radial acceleration?

OpenStudy (anonymous):

here tangential velocity will be zero ok meanns linear accn equals to acp radial accn

OpenStudy (anonymous):

solve this qnn as u knw a means linear accn

OpenStudy (australopithecus):

I got a = 0.4m/s^2 use the formula a = at + ar to find at so I have 0.4 =at + ar -6.25^2/0.2 = -31.25 0.4 + 31.25 = at this is probably incorrect

OpenStudy (anonymous):

as far i knw when we square v it becomes +Ve